OpenStudy (anonymous):

simplify the expression using (2x^3-11x^2+20x-12)/(x-2)

OpenStudy (anonymous):

Need help thanks!

OpenStudy (mathmate):

Factor the numberator: (2x^3-11x^2+20x-12)=(x-2)^2(2x-3) then cancel common factor (x-2) with the assumption that x is not equal to 2.

OpenStudy (anonymous):

okay then?

OpenStudy (mathmate):

Yes, unless you prefer to expand the remaining quadratic expression.

OpenStudy (anonymous):

okay so the answer is (x-2)

OpenStudy (mathmate):

Not quite: (x-2)^2(2x-3) / (x-2) = (x-2)(x-2)(2x-3) / (x-2) = (x-2)(2x-3) (with the requirement that x does not equal 2)

OpenStudy (anonymous):

so what does x equal?

OpenStudy (mathmate):

This problem is for simplifying, so no value will be given for x. Problems for solving (such as x=?) will provide you with equations (indicated by the presence of an equal sign), such as x^2-2x+1=0 can be solved to give x=1.

OpenStudy (anonymous):

Okay thank you, I struggle in math